Hagrid Company has a weighted average contribution margin per unit of $6.59. Fixed costs are $114,075 This yoga company sells three products, and anticipates the following sales volume next year. Yoga Videos Yoga Equipment Yoga Mats Total Units 13,000 units 15,000 units 12,000 units 40,000 units What is the breakeven quantity for Yoga Equipment next year?
Answer #1

Answer : Break Even Quantity ( Yoga Equipment ) = 6,491 Units

Break-Even Point = Fixed cost / Weighted Average Contribution Margin

Break-Even Point = 114,075 / 6.59 = 17,310 Units

Break even Point for Yoga Equipment = 17,310 *(15,000/40000) = 6,491 (Units)